# Hierarchical Retrieval with Evidence Curation for Open-Domain Financial Question Answering on Standardized Documents

### v2

@@ 5.6 Comparison of Commercial LLMs with Web Search
Method Numeric (Table) Numeric (Text) Textual Average
@@ 6 Conclusion
In our study, we use LLMs for query transformation, evidence curation, and answer generation, making our approach dependent on the performance of the LLMs. We utilize a relatively small LLM, Qwen 2.5 7B Yang et al. (2024), compared to commercial models.
@@ B.1 Categorization Details
Category Numeric (Text) Numeric (Table) Textual Total
@@
Category Numeric (Text) Numeric (Table) Textual Total
@@ Appendix E Generalization to Other Domains
To evaluate the effectiveness of our method beyond the financial domain, we assess retrieval performance on the LegalBench-RAG Pipitone and Alami (2024) dataset, a legal-domain benchmark where retrieval is particularly challenging due to domain-specific terminology and structured documents. Using a mini question set, we conduct retrieval over the full document corpus to simulate realistic conditions.

### v3

@@ 5.6 Comparison of Commercial LLMs with Web Search
Method
Numeric (Table)
Numeric (Text)
Textual
Average
@@ 6 Conclusion
In our study, we use LLMs for query transformation, evidence curation, and answer generation, making our approach dependent on the performance of the LLMs. We utilize a relatively small LLM, Qwen 2.5 7B yang2024qwen2, compared to commercial models.
@@ B.1 Categorization Details
Category
Numeric
Numeric
Textual Total
@@
Category
Numeric
Numeric
Textual Total
@@ Appendix E Generalization to Other Domains
To evaluate the effectiveness of our method beyond the financial domain, we assess retrieval performance on the LegalBench-RAG pipitone2024legalbench dataset, a legal-domain benchmark where retrieval is particularly challenging due to domain-specific terminology and structured documents. Using a mini question set, we conduct retrieval over the full document corpus to simulate realistic conditions.